OpenStudy (anonymous):

In how many arrangements can 3 boys and 4 girls stand in a row such that no two boys are together?

OpenStudy (anonymous):

@ganeshie8 @ParthKohli @AriPotta

Parth (parthkohli):

Number of ways that they can stand in - number of ways two boys can stand together. Note that the latter also covers the number of ways in which three boys can stand together.

Parth (parthkohli):

Oh, and also keep in mind that they can permute among themselves, like @ganeshie8 spotted the last time.

OpenStudy (anonymous):

I got 5040 at total number of ways they can stand together

OpenStudy (anonymous):

How would you calculate the number of ways two boys can stand together? I feel like I am messing up there

Parth (parthkohli):

Just like you did the last time.

Parth (parthkohli):

Oh, and yes, you have three boys, so you also need to see that you can choose any 2 out of 3 boys.

OpenStudy (anonymous):

See I messed up somewhere I got 5028...

ganeshie8 (ganeshie8):

4 girls can be arranged in 4! after that, 3 boys can be inserted in between the boys in 5C3 * 3! ways so in total we have 4! * 5C3*3! = 1440 ways

OpenStudy (anonymous):

720 4,320 3,600 4,000

Parth (parthkohli):

Why 5C3?

Parth (parthkohli):

\[7! - 2 \cdot \binom{3}{2}\cdot 6!\]

ganeshie8 (ganeshie8):

place girls in the blanks first 

Parth (parthkohli):

Oh, I see.

ganeshie8 (ganeshie8):

after that, we have 5 places for placing boys